OpenStudy (lizz123):

Medal One cell phone plan charges 20 dollars per month plus 0.15 cents per minute used. A second cell phone plan charges 35 dollars per month plus 0.10 cents per minute used. Find the number of minutes you must talk to have the same cost for both calling plans.

OpenStudy (michele_laino):

the cost for the first phone is: \[{c_1} = 20 + 0.15 \times m\] where m is the number of minutes. Whereas the cost for the second phone is: \[{c_2} = 35 + 0.10 \times m\]

OpenStudy (michele_laino):

now we have to know the number m of minute

OpenStudy (michele_laino):

minutes*

OpenStudy (lizz123):

\[c _{1}=20+.15\times1\] 20.15

OpenStudy (lizz123):

c1=20+.15*2 20.30

OpenStudy (michele_laino):

yes! nevertheless we have to find the value of m. In order to that we see that the cost c_1 is equal to the cost c_2 when the subsequent equation holds: \[\begin{gathered} {c_1} = {c_2} \hfill \\ 20 + 0.15 \times m = 35 + 0.10 \times m \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

please solve that equation for m

OpenStudy (lizz123):

Can I do a chart

OpenStudy (michele_laino):

if you want you can do a chart. We can get the answer if we solve that equation for m

OpenStudy (michele_laino):

\[\Large 20 + 0.15 \times m = 35 + 0.10 \times m\]

OpenStudy (lizz123):

C1=20+0.15*3= 20.45 C2=20+0.10*3= 20.330

OpenStudy (lizz123):

I mean 20.30

OpenStudy (lizz123):

So I have to find the same answer for the same time of minutes

OpenStudy (michele_laino):

ok! Please substitute m=100, what do you get?

OpenStudy (lizz123):

\[\huge c _{1}=20+.15 * 100 = 35\]

OpenStudy (michele_laino):

and c_2?

OpenStudy (lizz123):

\[\huge c _{2}= 35 +.10 * 100=45\]

OpenStudy (michele_laino):

ok! So they are different again. I think that the solution is m=300. Am I right?

OpenStudy (lizz123):

\[\huge c _{1}= 20 + .15*300= 65\] \[\huge c _{2}=35+.10*300=65\] Yes, m does equal 300.

OpenStudy (michele_laino):

perfect! So you can do your chart, using the subsequent values for m: 1, 20, 50, 100, 300 for example

OpenStudy (lizz123):

okay thanks

OpenStudy (michele_laino):

:)